# Simple homotopy types of even dimensional manifolds

### v2

@@ 4. Simple homotopy manifold sets
In this section we prove Theorem E about the characterisation of simple homotopy manifold sets. First we look at the analogous problem in the setting of CW complexes, which has a simpler answer. Then we consider the case of manifolds, which will rely on the results of Section 3.
@@ 4.1. CW complexes
Fix a CW complex $X$ and let $G=\pi_{1}(X)$. Our goal is to understand the set $\mathcal{C}^{h}_{s}(X)$ defined below.
@@
Let $\mathcal{C}^{h}_{s}(X)\coloneqq\left\{\text{CW complexes }Y\mid Y\simeq X\right\}/\simeq_{s}.$
We will need some auxiliary definitions.
@@
For $x\in\operatorname{Wh}(G)$ and $g\in\operatorname{hAut}(X)$ let $x^{g}=g_{*}(x)+\tau(g)$. Clearly $x^{\operatorname{Id}}=x$ and $x^{g\circ g^{\prime}}=(x^{g^{\prime}})^{g}$ (see Proposition 2.31), so this defines an action of $\operatorname{hAut}(X)$ on the set $\operatorname{Wh}(G)$.
With this notation we have the following theorem.
@@
$\widetilde{t}_{X}\colon\mathcal{C}^{h}_{s}(X)\to\operatorname{Wh}(G)/\operatorname{hAut}(X).$
@@
If $Y$ is a CW complex homotopy equivalent to $X$ and $Y\simeq_{s}Z$, then $t_{X}(Y)=t_{X}(Z)$.
@@
If $Y$ and $Z$ are CW complexes homotopy equivalent to $X$ and $t_{X}(Y)=t_{X}(Z)$, then $Y\simeq_{s}Z$.
@@
By Lemma 4.5, $t_{X}$ takes values in $\operatorname{Wh}(G)/\operatorname{hAut}(X)$ and by Lemma 4.6 it induces a well-defined map $\widetilde{t}_{X}$ on $\mathcal{C}^{h}_{s}(X)$. Lemmas 4.7 and 4.8 imply that $\widetilde{t}_{X}$ is injective and surjective, respectively. ∎
@@ 4.2. Manifolds
$\mathcal{M}^{h}_{s}(M)$$\coloneqq\left\{\text{closed $\operatorname{CAT}$ $n$-manifolds }N\mid N\simeq M\right\}/\simeq_{s}$
$\mathcal{M}^{\operatorname{hCob}}_{s}(M)$$\coloneqq\left\{\text{closed $\operatorname{CAT}$ $n$-manifolds }N\mid N\text{ is $h$-cobordant to }M\right\}/\simeq_{s}$
$\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)$$\coloneqq\left\{\text{closed $\operatorname{CAT}$ $n$-manifolds }N\mid N\simeq M\right\}/\langle\simeq_{s},\operatorname{hCob}\rangle$
where $\langle\simeq_{s},\operatorname{hCob}\rangle$ denotes the equivalence relation generated by simple homotopy equivalence and $h$-cobordism.
As before, $\operatorname{hAut}(M)$ acts on the set $\operatorname{Wh}(G,w)$, and a subset $t_{M}(N)\subseteq\operatorname{Wh}(G,w)$ is defined for every manifold $N$ that is homotopy equivalent to $M$. By Proposition 2.35 if $f\colon N\rightarrow M$ is a homotopy equivalence between manifolds, then $\tau(f)\in\mathcal{J}_{n}(G,w)$, so we can also define
@@
    $\widetilde{t}_{M}\colon\mathcal{M}^{h}_{s}(M)\to\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)\text{ and }\widetilde{t}^{\prime}_{M}\colon\mathcal{M}^{\operatorname{hCob}}_{s}(M)\to q(\mathcal{I}_{n}(G,w)),$
@@
    $\widetilde{u}_{M}\colon\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)\to\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))/\operatorname{hAut}(M).$
@@
    ${\mathcal{M}^{\operatorname{hCob}}_{s}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{t}^{\prime}_{M}}$${\mathcal{M}^{h}_{s}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{t}_{M}}$${\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{u}_{M}}$${q(\mathcal{I}_{n}(G,w))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))/\operatorname{hAut}(M)}$
@@
    If 1.1 are satisfied, then $\widetilde{t}^{\prime}_{M}$ is surjective, $\widetilde{u}_{M}$ is injective, and the top row is an exact sequence of pointed sets.
@@
    If 1.1 are satisfied, then, using the notation from Section 3.2, the subsets
    $\varrho^{-1}(\operatorname{Im}\sigma_{s})\subseteq\mathcal{J}_{n}(G,w)\text{ and }(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{s})\subseteq\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$
@@
    $\operatorname{Im}\widetilde{t}_{M}=(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{s})/\operatorname{hAut}(M)\text{ and }\operatorname{Im}\widetilde{u}_{M}=\varrho^{-1}(\operatorname{Im}\sigma_{s})/\operatorname{hAut}(M).$
This implies Theorem E (see also Theorem 4.16 below). Note that parts (a), (b) and (c) do not require 1.1 and so apply to smooth/PL 4-manifolds as well as topological 4-manifolds with arbitrary fundamental group.
We will use the following lemmas in the proof of Theorem 4.11.
@@
Let $N$ and $P$ be manifolds homotopy equivalent to $M$. If $N\simeq_{s}P$ or $N$ is $h$-cobordant to $P$, then $u_{M}(N)=u_{M}(P)$.
@@
If $N\simeq_{s}P$, then $t_{M}(N)=t_{M}(P)$ by Lemma 4.6, so $u_{M}(N)=\pi(t_{M}(N))=\pi(t_{M}(P))=u_{M}(P)$.
@@
Suppose that 1.1 are satisfied, and $N$ and $P$ are manifolds homotopy equivalent to $M$. If $u_{M}(N)=u_{M}(P)$, then there is a manifold $Q$ that is simple homotopy equivalent to $N$ and $h$-cobordant to $P$.
@@
Suppose that $N$ and $P$ are manifolds satisfying 1.1. If there is a manifold $Q$ that is simple homotopy equivalent to $N$ and $h$-cobordant to $P$, then there is a manifold $R$ that is $h$-cobordant to $N$ and simple homotopy equivalent to $P$.
@@
Suppose that $N$ and $P$ are manifolds satisfying 1.1. Then the following are equivalent.
  • (a)
@@
Now we are ready to begin the proof of Theorem 4.11.
@@
(a) If $g\in\operatorname{hAut}(M)$, then $\tau(g)\in\mathcal{J}_{n}(G,w)$ by Proposition 2.35. Moreover, $\pi_{1}(g)\circ w=w$, hence $g_{*}\colon\operatorname{Wh}(G,w)\rightarrow\operatorname{Wh}(G,w)$ is compatible with the involution, so if $x=-(-1)^{n}\overline{x}$, then $g_{*}(x)=-(-1)^{n}\overline{g_{*}(x)}$. Hence, if $x\in\mathcal{J}_{n}(G,w)$, then $x^{g}\in\mathcal{J}_{n}(G,w)$, i.e. $\operatorname{hAut}(M)$ acts on $\mathcal{J}_{n}(G,w)$.
@@
(b) If $f\colon N\rightarrow M$ is a homotopy equivalence for some $n$-manifold $N$, then $\tau(f)\in\mathcal{J}_{n}(G,w)$ by Proposition 2.35. Hence $t_{M}(N)\subseteq\mathcal{J}_{n}(G,w)$ for every manifold $N$ that is homotopy equivalent to $M$. Lemmas 4.5, 4.6, LABEL:, and 4.7 can be used again to show that $t_{M}$ induces well-defined injective maps $\widetilde{t}_{M}\colon\mathcal{M}^{h}_{s}(M)\to\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)$ and $\widetilde{t}^{\prime}_{M}\colon\mathcal{M}^{\operatorname{hCob}}_{s}(M)\to\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)$. If $N$ is $h$-cobordant to $M$ and $f\colon N\rightarrow M$ is a homotopy equivalence induced by an $h$-cobordism, then $\tau(f)\in\mathcal{I}_{n}(G,w)$ by Proposition 2.38, showing that $t_{M}(N)\in q(\mathcal{I}_{n}(G,w))$.
The analogue of Lemma 4.5 shows that $u_{M}(N)$ is an orbit of the action of $\operatorname{hAut}(M)$ on $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ (using that $\pi(\tau(g\circ f))=\pi(\tau(f)^{g})=\pi(\tau(f))^{g}$ for every homotopy equivalence $f\colon N\to M$ and $g\in\operatorname{hAut}(M)$). Lemma 4.12 shows that $u_{M}$ induces a well-defined map on $\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)$.
@@
For the exactness of the top row, assume that $N$ is equivalent to $M$ under the equivalence relation generated by simple homotopy equivalence and $h$-cobordism. By Proposition 4.15 there is a manifold $P$ that is simple homotopy equivalent to $N$ and $h$-cobordant to $M$. This represents an element of $\mathcal{M}^{\operatorname{hCob}}_{s}(M)$, which is mapped to the equivalence class of $N$ in $\mathcal{M}^{h}_{s}(M)$.
@@
(e) First consider $\varrho^{-1}(\operatorname{Im}\sigma_{s})$, which is equal to $\operatorname{Im}\widehat{\tau}$ by Corollary 3.9. This is $\operatorname{hAut}(M)$-invariant, because for any $[N,f]\in\mathcal{S}^{h}(M)$ and $g\in\operatorname{hAut}(M)$ we have $\pi(\tau(f))^{g}=\pi(\tau(g\circ f))=\widehat{\tau}([N,g\circ f])$. Since the action of $\operatorname{hAut}(M)$ on $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is induced by its action on $\mathcal{J}_{n}(G,w)$, and $\varrho^{-1}(\operatorname{Im}\sigma_{s})$ is $\operatorname{hAut}(M)$-invariant, $\pi^{-1}(\varrho^{-1}(\operatorname{Im}\sigma_{s}))$ is $\operatorname{hAut}(M)$-invariant too.
@@
Now suppose that $[N]\in\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)$. Then there is a homotopy equivalence $f\colon N\rightarrow M$ and $\widetilde{u}_{M}([N])$ is the orbit of $\pi(\tau(f))=\widehat{\tau}([N,f])\in\operatorname{Im}\widehat{\tau}=\varrho^{-1}(\operatorname{Im}\sigma_{s})$. This shows that $\operatorname{Im}\widetilde{u}_{M}\subseteq\varrho^{-1}(\operatorname{Im}\sigma_{s})/\operatorname{hAut}(M)$. On the other hand, if $x\in\varrho^{-1}(\operatorname{Im}\sigma_{s})=\operatorname{Im}\widehat{\tau}$, then $x=\pi(\tau(f))$ for some homotopy equivalence $f\colon N\rightarrow M$, and then $[N]\in\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)$ and $\widetilde{u}_{M}([N])$ is the orbit of $x$. Therefore $\operatorname{Im}\widetilde{u}_{M}\supseteq\varrho^{-1}(\operatorname{Im}\sigma_{s})/\operatorname{hAut}(M)$.
@@
By the commutativity of the diagram in part (c), $\operatorname{Im}\widetilde{t}_{M}$ is contained in the inverse image of $\operatorname{Im}\widetilde{u}_{M}=\varrho^{-1}(\operatorname{Im}\sigma_{s})/\operatorname{hAut}(M)$, which is $\pi^{-1}(\varrho^{-1}(\operatorname{Im}\sigma_{s}))/\operatorname{hAut}(M)$. In the other direction, if $x\in\mathcal{J}_{n}(G,w)$ and $\pi(x)\in\varrho^{-1}(\operatorname{Im}\sigma_{s})=\operatorname{Im}\widehat{\tau}$, then there exists a manifold $N_{1}$ and a homotopy equivalence $f_{1}\colon N_{1}\to M$ with $\pi(\tau(f_{1}))=\pi(x)$. It follows that $x=\tau(f_{1})+y$ for some $y\in\mathcal{I}_{n}(G,w)$. Use $(f_{1})_{*}$ to identify $\pi_{1}(N_{1})$ with $G=\pi_{1}(M)$ (since $f_{1}$ is a homotopy equivalence, the orientation character of $N$ is also $w$). By Corollary 3.2 there exists a manifold $N_{2}$ and a homotopy equivalence $f_{2}\colon N_{2}\to N_{1}$ with $\tau(f_{2})=y$. By Proposition 2.31 we have $\tau(f_{1}\circ f_{2})=\tau(f_{1})+(f_{1})_{*}(\tau(f_{2}))=\tau(f_{1})+\tau(f_{2})=\tau(f_{1})+y=x$ (where $(f_{1})_{*}=\operatorname{Id}$ because we used $f_{1}$ to identify $\pi_{1}(N_{1})$ with $G$). So for $[N_{2}]\in\mathcal{M}^{h}_{s}(M)$ we get that $\widetilde{t}_{M}([N_{2}])$ is the orbit of $x$, showing that $\operatorname{Im}\widetilde{t}_{M}\supseteq(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{s})/\operatorname{hAut}(M)$. ∎
@@
If 1.1 are satisfied, then there are bijections
$\widetilde{t}^{\prime}_{M}$ $\colon\mathcal{M}^{\operatorname{hCob}}_{s}(M)\xrightarrow{\cong}q(\mathcal{I}_{n}(G,w))$
$\widetilde{t}_{M}$ $\colon\mathcal{M}^{h}_{s}(M)\xrightarrow{\cong}(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{s})/\operatorname{hAut}(M)$
$\widetilde{u}_{M}$ $\colon\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)\xrightarrow{\cong}\varrho^{-1}(\operatorname{Im}\sigma_{s})/\operatorname{hAut}(M).$
@@
If 1.1 are satisfied and $\psi$ is surjective, then there are bijections $\mathcal{M}^{h}_{s}(M)\cong\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)$ and $\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)\cong\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))/\operatorname{hAut}(M)$.

### v3

@@ 4. Simple homotopy manifold sets
In this section we prove Theorem B about the characterisation of simple homotopy manifold sets. First we look at the analogous problem in the setting of CW complexes, which has a simpler answer. Then we consider the case of manifolds, which will rely on the results of Section 3.
@@ 4.1. CW complexes
Fix a (finite, connected) CW complex $X$ and let $G=\pi_{1}(X)$. Our goal is to understand the set $\mathcal{C}^{\mathrm{h}}_{\mathrm{s}}(X)$ defined below.
@@
Let $\mathcal{C}^{\mathrm{h}}_{\mathrm{s}}(X):=\left\{\text{CW complexes }Y\mid Y\simeq X\right\}/\simeq_{\mathrm{s}}.$
Note that the representatives of elements of $\mathcal{C}^{\mathrm{h}}_{\mathrm{s}}(X)$ consist only of CW complexes $Y$, and do not include chosen homotopy equivalences $Y\rightarrow X$.
We now need some auxiliary definitions.
@@
For $x\in\operatorname{Wh}(G)$ and $g\in\operatorname{hAut}(X)$ let $x^{g}=g_{*}(x)+\tau(g)$.
Clearly $x^{\operatorname{Id}}=x$ and $x^{g\circ g^{\prime}}=(x^{g^{\prime}})^{g}$ (see Proposition 2.31), so this defines an action of $\operatorname{hAut}(X)$ on the set $\operatorname{Wh}(G)$. With this notation we have the following theorem.
@@
$\widetilde{t}_{X}\colon\mathcal{C}^{\mathrm{h}}_{\mathrm{s}}(X)\to\operatorname{Wh}(G)/\operatorname{hAut}(X).$
@@
If $Y$ is a CW complex homotopy equivalent to $X$ and $Y\simeq_{\mathrm{s}}Z$, then $t_{X}(Y)=t_{X}(Z)$.
@@
If $Y$ and $Z$ are CW complexes homotopy equivalent to $X$ and $t_{X}(Y)=t_{X}(Z)$, then $Y\simeq_{\mathrm{s}}Z$.
@@
By Lemma 4.5, $t_{X}$ takes values in $\operatorname{Wh}(G)/\operatorname{hAut}(X)$ and by Lemma 4.6 it induces a well-defined map $\widetilde{t}_{X}$ on $\mathcal{C}^{\mathrm{h}}_{\mathrm{s}}(X)$. Lemmas 4.7 and 4.8 imply that $\widetilde{t}_{X}$ is injective and surjective, respectively. ∎
@@ 4.2. Manifolds
$\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)$ $=\left\{\text{closed $\operatorname{CAT}$ $n$-manifolds }N\mid N\simeq M\right\}/\simeq_{\mathrm{s}}$
$\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)$ $=\left\{\text{closed $\operatorname{CAT}$ $n$-manifolds }N\mid N\text{ is $h$-cobordant to }M\right\}/\simeq_{\mathrm{s}}$
$\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)$ $=\left\{\text{closed $\operatorname{CAT}$ $n$-manifolds }N\mid N\simeq M\right\}/\langle\simeq_{\mathrm{s}},\operatorname{hCob}\rangle$
where $\langle\simeq_{\mathrm{s}},\operatorname{hCob}\rangle$ denotes the equivalence relation generated by simple homotopy equivalence and $h$-cobordism.
As in Section 4.1, $\operatorname{hAut}(M)$ acts on the set $\operatorname{Wh}(G,w)$, and a subset $t_{M}(N)\subseteq\operatorname{Wh}(G,w)$ is defined for every manifold $N$ that is homotopy equivalent to $M$. Let $\pi\colon\mathcal{J}_{n}(G,w)\rightarrow\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ denote the quotient map. By Proposition 2.35 if $f\colon N\rightarrow M$ is a homotopy equivalence between manifolds, then $\tau(f)\in\mathcal{J}_{n}(G,w)$, so we can also define
@@
    $\widetilde{t}_{M}\colon\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)\to\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)\text{ and }\widetilde{t}^{\prime}_{M}\colon\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)\to q(\mathcal{I}_{n}(G,w)),$
@@
    $\widetilde{u}_{M}\colon\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)\to\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))/\operatorname{hAut}(M).$
@@
    ${\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{t}^{\prime}_{M}}$${\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{t}_{M}}$${\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{u}_{M}}$${q(\mathcal{I}_{n}(G,w))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))/\operatorname{hAut}(M)}$
@@
    If 1.1 is satisfied, then $\widetilde{t}^{\prime}_{M}$ is surjective, $\widetilde{u}_{M}$ is injective, and the top row is an exact sequence of pointed sets.
@@
    If 1.1 is satisfied, then, using the notation from Section 3.2, the subsets
    $\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})\subseteq\mathcal{J}_{n}(G,w)\text{ and }(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})\subseteq\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$
@@
    $\operatorname{Im}\widetilde{t}_{M}=(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})/\operatorname{hAut}(M)\text{ and }\operatorname{Im}\widetilde{u}_{M}=\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})/\operatorname{hAut}(M).$
This implies Theorem B (see also Theorem 4.16 below). Parts (a), (b) and (c) do not require 1.1 and so apply to smooth/PL 4-manifolds as well as topological 4-manifolds with arbitrary fundamental group. Parts (d) and (e) rely on results from Section 3, and so they require 1.1.
We use the following lemmas in the proof of Theorem 4.11.
@@
Let $N$ and $P$ be manifolds homotopy equivalent to $M$. If $N\simeq_{\mathrm{s}}P$ or $N$ is $h$-cobordant to $P$, then $u_{M}(N)=u_{M}(P)$.
@@
If $N\simeq_{\mathrm{s}}P$, then $t_{M}(N)=t_{M}(P)$ by Lemma 4.6, so $u_{M}(N)=\pi(t_{M}(N))=\pi(t_{M}(P))=u_{M}(P)$.
@@
Suppose that 1.1 is satisfied, and $N$ and $P$ are $\operatorname{CAT}$ $n$-manifolds homotopy equivalent to $M$. If $u_{M}(N)=u_{M}(P)$, then there is a manifold $Q$ that is simple homotopy equivalent to $N$ and $h$-cobordant to $P$.
@@
Suppose that $N$ and $P$ are CATCAT $n$-manifolds with $\operatorname{CAT}$ as in 1.1. If there is a manifold $Q$ that is simple homotopy equivalent to $N$ and $h$-cobordant to $P$, then there is a manifold $R$ that is $h$-cobordant to $N$ and simple homotopy equivalent to $P$.
@@
Suppose that $N$ and $P$ are $\operatorname{CAT}$ $n$-manifolds with $\operatorname{CAT}$ as in 1.1. Then the following are equivalent.
  1. (a)
@@
We are now ready to begin the proof of Theorem 4.11.
@@
(a) If $g\in\operatorname{hAut}(M)$, then $\tau(g)\in\mathcal{J}_{n}(G,w)$ by Proposition 2.35. Moreover, $H_{n}(M;\mathbb{Z}^{w\circ\pi_{1}(g)})\cong H_{n}(M;\mathbb{Z}^{w})\cong\mathbb{Z}$, so $w\circ\pi_{1}(g)=w$. Thus $g_{*}\colon\operatorname{Wh}(G,w)\rightarrow\operatorname{Wh}(G,w)$ is compatible with the involution, so if $x=-(-1)^{n}\overline{x}$, then $g_{*}(x)=-(-1)^{n}\overline{g_{*}(x)}$. Hence, if $x\in\mathcal{J}_{n}(G,w)$, then $x^{g}\in\mathcal{J}_{n}(G,w)$, i.e. $\operatorname{hAut}(M)$ acts on $\mathcal{J}_{n}(G,w)$.
@@
(b) If $f\colon N\rightarrow M$ is a homotopy equivalence for some $n$-manifold NN, then $\tau(f)\in\mathcal{J}_{n}(G,w)$ by Proposition 2.35. Hence $t_{M}(N)\subseteq\mathcal{J}_{n}(G,w)$ for every manifold $N$ that is homotopy equivalent to $M$. Lemmas 4.5, 4.6, LABEL:, and 4.7 can be used again to show that $t_{M}$ induces well-defined injective maps $\widetilde{t}_{M}\colon\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)\to\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)$ and $\widetilde{t}^{\prime}_{M}\colon\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)\to\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)$. If $N$ is $h$-cobordant to $M$ and $f\colon N\rightarrow M$ is a homotopy equivalence induced by an $h$-cobordism, then $\tau(f)\in\mathcal{I}_{n}(G,w)$ by Proposition 2.38, showing that $t_{M}(N)\in q(\mathcal{I}_{n}(G,w))$.
The analogue of Lemma 4.5 shows that $u_{M}(N)$ is an orbit of the action of $\operatorname{hAut}(M)$ on $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ (using that $\pi(\tau(g\circ f))=\pi(\tau(f)^{g})=\pi(\tau(f))^{g}$ for every homotopy equivalence $f\colon N\to M$ and $g\in\operatorname{hAut}(M)$). Lemma 4.12 shows that $u_{M}$ induces a well-defined map on $\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)$.
@@
For the exactness of the top row, assume that $N$ is equivalent to $M$ under the equivalence relation generated by simple homotopy equivalence and $h$-cobordism. By Proposition 4.15 there is a manifold $P$ that is simple homotopy equivalent to $N$ and $h$-cobordant to $M$. This represents an element of $\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)$, which is mapped to the equivalence class of $N$ in $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)$.
@@
(e) First consider $\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})$, which is equal to $\operatorname{Im}\widehat{\tau}$ by Corollary 3.11. This is $\operatorname{hAut}(M)$-invariant, because for any $N,f]\in\mathcal{S}^{\mathrm{h}}(M)$ and $g\in\operatorname{hAut}(M)$ we have $\pi(\tau(f))^{g}=\pi(\tau(g\circ f))=\widehat{\tau}([N,g\circ f])$. Since the action of $\operatorname{hAut}(M)$ on $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is induced by its action on $\mathcal{J}_{n}(G,w)$, and $\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})$ is $\operatorname{hAut}(M)$-invariant, $\pi^{-1}(\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}}))$ is $\operatorname{hAut}(M)$-invariant too.
@@
Now suppose that $N]\in\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)$. Then there is a homotopy equivalence $f\colon N\rightarrow M$ and $\widetilde{u}_{M}([N])$ is the orbit of $\pi(\tau(f))=\widehat{\tau}([N,f])\in\operatorname{Im}\widehat{\tau}=\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})$. This shows that $\operatorname{Im}\widetilde{u}_{M}\subseteq\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})/\operatorname{hAut}(M)$. On the other hand, if $x\in\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})=\operatorname{Im}\widehat{\tau}$, then $x=\pi(\tau(f))$ for some homotopy equivalence $f\colon N\rightarrow M$, and then $N]\in\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)$ and $\widetilde{u}_{M}([N])$ is the orbit of $x$. Therefore $\operatorname{Im}\widetilde{u}_{M}\supseteq\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})/\operatorname{hAut}(M)$.
@@
By the commutativity of the diagram in part (c), $\operatorname{Im}\widetilde{t}_{M}$ is contained in the inverse image of $\operatorname{Im}\widetilde{u}_{M}=\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})/\operatorname{hAut}(M)$, which is $\pi^{-1}(\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}}))/\operatorname{hAut}(M)$. In the other direction, if $x\in\mathcal{J}_{n}(G,w)$ and $\pi(x)\in\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})=\operatorname{Im}\widehat{\tau}$, then there exists a manifold $N_{1}$ and a homotopy equivalence $f_{1}\colon N_{1}\to M$ with $\pi(\tau(f_{1}))=\pi(x)$. It follows that $x=\tau(f_{1})+y$ for some $y\in\mathcal{I}_{n}(G,w)$. Use $(f_{1})_{*}$ to identify $\pi_{1}(N_{1})$ with $G=\pi_{1}(M)$ (since $f_{1}$ is a homotopy equivalence, the orientation character of $N$ is also $w$). By Corollary 3.2 there exists a manifold $N_{2}$ and a homotopy equivalence $f_{2}\colon N_{2}\to N_{1}$ with $\tau(f_{2})=y$. By Proposition 2.31 we have $\tau(f_{1}\circ f_{2})=\tau(f_{1})+(f_{1})_{*}(\tau(f_{2}))=\tau(f_{1})+\tau(f_{2})=\tau(f_{1})+y=x$ (where $(f_{1})_{*}=\operatorname{Id}$ because we used $f_{1}$ to identify $\pi_{1}(N_{1})$ with $G$). So for $N_{2}]\in\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)$ we get that $\widetilde{t}_{M}([N_{2}])$ is the orbit of $x$, showing that $\operatorname{Im}\widetilde{t}_{M}\supseteq(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})/\operatorname{hAut}(M)$. ∎
@@
If 1.1 is satisfied, then there are bijections
$\widetilde{t}^{\prime}_{M}$ $\colon\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)\xrightarrow{\cong}q(\mathcal{I}_{n}(G,w))$
$\widetilde{t}_{M}$ $\colon\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)\xrightarrow{\cong}(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})/\operatorname{hAut}(M)$
$\widetilde{u}_{M}$ $\colon\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)\xrightarrow{\cong}\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})/\operatorname{hAut}(M).$
@@
If 1.1 is satisfied and the map $\psi$ is surjective, then there are bijections $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)\cong\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)$ and $\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)\cong\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))/\operatorname{hAut}(M)$.